OpenStudy (anonymous):

Suppose f(x)=x^9+5 and g(x)=sqrt x . Then what is f(g(x))and g(f(x))

OpenStudy (anonymous):

f(g(x))=(sqrtx)^9+5

OpenStudy (anonymous):

g(f(x))=sqrt(x^9+5)

OpenStudy (anonymous):

just substitute one function into another

OpenStudy (anonymous):

i need help with the f(g(x)) do i have to simplify the \[\sqrt{x}\]^9

OpenStudy (anonymous):

\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[x ^{\frac{ 1 }{ 2 }*9}=x ^{\frac{ 9 }{ 2 }}\]

OpenStudy (anonymous):

thank you